Date Price before Dividend Dividend 2000 2001 2002 2003 2004 2005 2006 2007 2008 2009 2010 2011 2012 2013 2014 2015 742.5 758.8 742.9 697.7 744.9 693 735.7 692 768.7 767 12.6 11.1 15.6 18.2 18 19.7 19.2 21.1 20.5 21.5 22.4 23 23.3 24.1 25 784.5 758 796.7 823.2(1). Calculate the gross annual return on this stock and report the number in a separate column.

(2). Calculate the sample mean for the gross return.

(3). Calculate the sample standard deviation for the gross return.

(4). Assume the risk-free rate is 0.02. Estimate the risk premium.

(5). Estimate the sharp ratio.
Answer #1

1) Gross Annual Return

Gross Annual Return = Dividend/Price * 100

Year Price Dividend Annual Return
2000 742.5 12.6 1.70%
2001 758.8 11.1 1.46%
2002 742.90 15.6 2.10%
2003 697.7 18.2 2.61%
2004 744.9 18 2.42%
2005 693 19.7 2.84%
2006 735.7 19.2 2.61%
2007 692 21.1 3.05%
2008 768.7 20.5 2.67%
2009 767 21.5 2.80%
2010 777.2 22 2.83%
2011 784.5 22.4 2.85%
2012 758 23 3.03%
2013 771.8 23.3 3.02%
2014 796.7 24.1 3.02%
2015 823.2 25 3.04%

2.

Sample Mean of Gross Return

= (1.7+1.46+2.10+2.61+2.42+2.84+2.61+3.05+2.67+2.80+2.83+2.85+3.03+3.02+3.02+3.04)/16

= 42.05/16

= 2.63%

3. Sample Standard Deviation

= (Annual Return - Mean Return)^2

= 3.52

= 3.52/16

= 0.22

Under root of 0.22

= 0.47%

Standard deviation = 0.47%

5. Sharpe ratio = (Mean portfolio return − Risk-free rate)/Standard deviation of portfolio return

= (0.0263 - 0.02)/0.0047

= 1.34

4. Risk Premium

The estimated stock return minus the risk free rate equals the estimated risk premium.

= Mean - Risk free rate

= 0.0263 - 0.02

= 0.63% or 0.0063